OpenStudy (anonymous):

please help ...Write the equation of the line that is parallel to the line 4x − 3y = −12 and passes through the point (−3, 4)

OpenStudy (anonymous):

answer choices are: y = 4/3x + 8 y = 4/3x + 3 y = -3/4 +8 y = -3/4x +3

OpenStudy (anonymous):

for it to be paralell, we don,t really care about the Right Hand Side (RHS) aka -12 so we just bring that line to the origin which gives: 4x-3y=0, then solving for y we get: y = (4/3)x + c (c is any number dependint on what the intercept is), now just plug in your point (-3,4) we get 4 = (4/3)*(-3) + c ----> 4 = -4 + c ----> c = 8 hence your final awnser is: y = (4/3)x + 8 have a nice day! :)

OpenStudy (anonymous):

wow your smart thanks